calculate the aggregate expenditure and multiplier when C= 300+0.14Y, I= 700, G = 5000, TX= 3+ 0.0Y and TR= 90


1
Expert's answer
2021-09-07T13:18:23-0400

"C=300+0.14Y"

"I=700"

"G=5000"

"NX= 3+0.0Y"

"Y= C+I+G"

"Y=300+0.14Y+700+5000"

"Y=6976.74"

"AE= C+I+G+NX= Y+NX"

"AE= 6976.74+ 3+0.0(6976.74)= 6979.7436"


"MPC= \frac { 300+0.14(6976.74)}{6976.74}=0.183"

Multiplier"= \frac {1}{1-0.183}=1.2"